OpenStudy (anonymous):

Divide. 4u^2 - 6u/ 2u A. 4u - 6/ 2 B. 4u - 6 C. 2u - 3/ 2 D. 2u - 3

OpenStudy (ajprincess):

is ur question this? \[\frac{4u^2-6u}{2u}\]

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

???????

OpenStudy (anonymous):

can you just tell me you dont have to give big explanation

OpenStudy (ajprincess):

k:) \[\frac{4u^2-6u}{2u}\] \[=\frac{4u^2}{2u}-\frac{6u}{2u}\] \[=\frac{2u*2u}{2u}-\frac{3*2u}{2u}\]\[=\frac{2u(2u-3)}{2u}\] Cancel any common factorss nd simplify.:) Does tht help? @Tcookie009

OpenStudy (anonymous):

wich one is it

OpenStudy (ajprincess):

what is common to both numerator and denominator?

OpenStudy (anonymous):

can you just tell me and ill give you three medals

OpenStudy (ajprincess):

I am really sorry i am afraid i cannot giv u the direct answer and I am nt trying to help u for the medals.:)

OpenStudy (anonymous):

wich one are the wrrong ones

OpenStudy (ajprincess):

can u pls tell me what u see in both numerator and denominator of this fraction? \[\frac{2u(2u-3)}{2u}\]